How many sides does a Decagon have?

A

7

B

8

C

9

D

10

 

Grade:12

2 Answers

Vikas TU
2 years ago
Dear student 
Decagon have 10 sides , and each angle is 144 degreee. 
Hope this will help 
Feel free to ask querries. 
Good Luck 
Chers 
ankit singh
2 years ago
Explanation: We know that to calculate number of diagonals in octagon is Number of diagonals is= n × (n -3)/2 27= n × (n -3)/2 n (n-3) = 54 n2– 3n = 54 n2– 3n-54 = 0 (n + 6) (n – 9)=0 n=-6 or n=9 So that we are calculating the sides it should be positive, therefore sides of polygon has 27 diagonals is 9